Solve, show all work
If the first and third of three consecutive odd integers are added, the result is 69 less than five times the second integer. Find all three integers.

Answer by robertb(5830)

x, x+2, x+4
==> x+(x+4) = 5(x+2)-69
==> 2x + 4 = 5x + 10 - 69
==> 2x + 4 = 5x - 59
==> 63 = 3x
==> x = 21, the 1st odd integer.
23, the second,
25, the last odd integer.